Find the x- and y-intercepts of 7x – 2y = –5 and use these intercepts to graph...
I got (-5/7, -5/2) not sure if they are correct and cannot graph

Answer by nerdybill(7384)

.
The x and y intercepts are the points at which the line crosses the x-axis and y-axis respectively. Therefore, for the y-intercept the FIRST coordinate will be zero whereas for the x-axis the SECOND coordinate will be zero. In other words, to find the y-intercept, set 'x' equal to zero and solve for y. And for the x-intercept, set 'y' equal to zero and solve for x.
.
Finding the y-intercept:
7x – 2y = –5
Now, set 'x' = 0 and solve for y:
7(0) – 2y = –5
– 2y = –5
y = 5/2
So, the y-intercept is at (0, 5/2)
Plot this point on graph paper.
.
Finding the x-intercept:
7x – 2y = –5
Now, set 'y' = 0 and solve for x:
7x – 2(0) = –5
7x = –5
x = -5/7
So, the x-intercept is at (-5/7, 0)
Plot this point on graph paper.
.
Finally, to complete the graph, connect the two points and now you're done